# Planning and Learning in Risk-Aware Restless Multi-Arm Bandit Problem

### v1

@@ 5 Numerical Analysis
Figure 3: The plot shows $\mathcal{R}(k)$ of Risk-aware Whittle index policy on the top and $\mathcal{R}(k)/K$ on the bottom for three different setups when the utility function of $(\alpha,\tau)=(1,0.5)$. For all these experiments, $T=5$ and $M=1$. These plots are averaged over $100$ sample paths. The transition model for figures $(a)\&(b)$ and $(c)\&(d)$ are according to $\mathcal{P}$, for figures $(e)\&(f)$ is according to model 1 of Le, (2016).
@@
We assume the Markov chain for each treatment and arm follows model 1 from the article, with transition probabilities sampled from the specified range. For the learning problem, the true transition probabilities are unknown, so we set a Dirichlet prior for Thompson sampling.
@@
At the start of each episode, the risk-aware Whittle index policy is computed based on estimated parameters. During the episode, the learner observes 10 batches of finite-horizon rounds. State-action-state observations for each arm are counted, and Dirichlet distributions are constructed over the unknown parameters. These parameters are then updated by sampling from the distributions.
@@
In Figure 3, we illustrate $\mathcal{R}{(k)}$ for a learner implementing RAWIP-TS, compared to an oracle with full knowledge of the true parameters who implements the true risk-aware Whittle index policy. The results suggest the learning mechanism is effective, with regret growing sublinearly as $\mathcal{R}{(k)}/K$ converges to 0. Although the risk-aware Whittle index policy is not optimal, leading to possible negative regret (as in cases (b) and (c)), the learner’s policy eventually converges to the oracle’s policy.

### v2

@@ 5 NUMERICAL ANALYSIS
Figure 3: The plot shows $\mathcal{R}(k)$ of RAWIP on the left and $\mathcal{R}(k)/K$ on the right for three different setups when the utility function is $\alpha=1$. For all these experiments, $T=5$ and $M=1$, $\tau=0.5$. These plots are averaged over $100$ sample paths. The transition model for figures $(a)\&(b)$ is according to model 3 of Le (2016) and $(c)\&(d)$ and $(e)\&(f)$ are according to $\mathcal{P}$.
The table illustrates the average performance of our algorithm (computed across multiple problem setups and iterations) as parameters change. For utility functions 2 and 3, the results show that increasing the order—moving further from the linear case—highlights the superior performance of the risk-aware policy. This is expected, as risk-aware and risk-neutral policies perform identically under a linear utility function.
Learning: This section outlines the experiments conducted for 2.4. Since computing the optimal policy is infeasible for large numbers of arms, even with estimated or true parameters, we use the risk-aware Whittle index policy (Section 3.4, which we refer to as RAWIP) as a proxy. We refer to the algorithm as Risk-Aware Whittle Index Policy using Thompson Sampling (RAWIP-TS). We consider three models: one for patient scheduling and two for machine maintenance scheduling.
For patient scheduling problem, each arm represents a patient with Advanced Breast Cancer,  Le (2016), where states represent disease progression (DEAD: 0, PROGRESSING: 1, STABLE: 2). Patients are treated with either Capecitabine or Lapatinib+Capecitabine. Due to limited hospital resources, the problem is modeled as a restless bandit problem with a fixed budget. The rewards for each state are: $r(\text{DEAD})=0$, $r(\text{PROGRESSING})=1/(2T)$, $r(\text{STABLE})=1/(T)$. We use model 3 described in Le (2016) as the state dynamics. According to Table 1 of the paper, there’s a range for for each transition probabilities. We first choose a value for each transition probability uniformly random from the given range and assuring the sum of transitions from a given state is $1$, and then run the experiments.
Remark 9.
The randomly generated Markov dynamics from Le (2016) may not satisfy our assumptions mentioned in Theorem 5 for indexability. However, the problem instances are indexable. This showcases that our approach works for any indexable RMABs and the assumptions mentioned in Theorem 5 are only a set of sufficient conditions.
For machine maintenance problem, each arm represents a machine with exactly the same setup as the one described in the planning part. However, the true values of $p^{i}$ are unknown to the learner. We assume $p^{i}\in[0.1/|\mathcal{X}|,1/|\mathcal{X}|]$ for arm $i$, and set the initial estimate $\hat{p}^{i}_{0}$ to the median value $0.5/|\mathcal{X}|$.
@@
For the learning problem, the true transition probabilities are unknown, so we set a Dirichlet prior for Thompson sampling.
@@
At the start of each episode, the RAWIP is computed based on estimated parameters. During the episode, the learner observes 10 batches of finite-horizon rounds. State-action-state observations for each arm are counted, and Dirichlet distributions are constructed over the unknown parameters. These parameters are then updated by sampling from the distributions.
@@
In Figure 3, we illustrate $\mathcal{R}{(k)}$ for a learner implementing RAWIP-TS, compared to an oracle with full knowledge of the true parameters who implements the true RAWIP. The results suggest the learning mechanism is effective, with regret growing sublinearly as $\mathcal{R}{(k)}/K$ converges to 0. Note as the RAWIP is generally suboptimal, the RAWIP-TS may outperform it and hence, the slope of cumulative regret becomes negative (as in case (a)). However, the learner’s policy eventually converges to the RAWIP (as shown in (b)).